On January 1, 2016, Bennett Corporation had 20,000 shares of common shares outstanding. Bennett sold another 2,600 shares on July 1, 2016, and reacquired 600 shares on November 1, 2016. Bennett earned $337,600 net income and has 15,000 shares of $10 par value, 6% cumulative preferred stock on which the last dividends were declared in 2012. Select the year’s basic earnings per share.

$15.92
$15.65
$15.50
$15.08
Answer #1

Preferred dividend = 15000*10*6% = 9000

Weighted average share outstanding = (20000*6/12+22600*4/12+22000*2/12) = 21200 Shares

Earning per share = (337600-9000)/21200 = 15.50

So answer is c) $15.50